# Low-budget Black-box Optimization Algorithms Evaluated on BBOB and OpenAI Gym

### v1

@@ III Benchmark Problems
Winning percentage rates are evaluated at fixed budgets of 25, 50, 100, 200, and 400 function evaluations for both settings (neural factor $1$ and $3$), while bigger budgets are also considered for the bigger networks (see Appendices D and E).
@@ IV Results of the Empirical Comparison
Fig. 2 shows that both versions of CMA – defaultcma and cmafmin2 – perform better than all other solvers when a larger budget of function evaluations is available.
@@
We add the following algorithms to our comparison: QORandomSearch [67], MetaTuneRecentering [68], and MetaRecentering [69] are variants of random search that are fully parallel and reduce redundancies compared to random search. We also add NGOpt16RL and SpecialRL from Nevergrad: NGOpt16RL is Nevergrad’s wizard NGOpt optimized specifically for RL problems; SpecialRL runs the base NGOpt16RL for half of the budget and then uses test-based population size adaptation (TBPSA) [72] in the second half. Moreover, since the OpenAI Gym suite contains problems that are unbounded with optima at a small scale, we add algorithms that can adjust the scale of the search. GeneticDE and RotatedTwoPointsDE, for example, are designed to extrapolate the scale from some variables to others and find the right scale before running the classic DE algorithm. MetaTuneRecentering performs a DoE entirely designed to choose the right scale depending on the relationship between budget and dimension. MultiCMA is a tentative robustification of CMA that runs three copies of CMA. We also consider a (1+1) evolutionary algorithm [52], which is a simple and fast state-of-the-art heuristic for adjusting the scale, Diagonal CMA [27], which uses a diagonal covariance matrix, and its scaled version, Scaled Diagonal CMA [5]. Finally, we add to our portfolio Lamcts-Turbo [18], an improved version of vanilla LA-MCTS coupled with Turbo to draw new samples in the selected node of the Monte Carlo search tree at each iteration. For a complete list of all algorithms compared, see Appendix F.
@@
Figure 4: Same results as reported in Fig. 3, but aggregated comparison as provided by Nevergrad [5]. Row A column B shows the frequency at which method A outperformed method B for the given budget. 9 distinct problems per budget. We include only problems for which the dimension is $D<50$. Methods are ranked per average winning rate; best methods are listed first. Note that winning rates are all very close to each other: only PSO is significantly better. Fig. 5 presents similar experiments but with bigger neural nets. Fig. 6 extends the present results to budgets 800, 1600, and 3200.
@@
Figure 5: Same as Fig. 4, but with bigger nets (neural factor 3 in Nevergrad’s benchmark scaling). 11 distinct problems per budget. We truncated at dimension $\leq 264$. Dimension ranges from 24 to 264 instead of 8 to 40 in Fig. 4. Due to the computational cost, it was not possible to finish the runs for SMAC. Fig. 6 extends the present results to budgets 800, 1600, and 3200.
For an in-depth comparison, we also present aggregate plots based on average winning rates (e.g., Fig. 4), where we observe good results for the BO-based methods: BO is the most powerful solver for the lowest budget, while AX turns out to be the best algorithm when aggregating results across all budgets (Fig. 4f). This means that vanilla BO is actually good for fast low-precision approximations on difficult problems, whereas AX can be recommended when optimizing low-dimensional problems with low to medium budgets. Furthermore, Fig. 4 confirms, on the one hand, the good performance of NGOpt16RL, PSO, and HyperOpt for benchmarks with tiny neural networks, and, on the other hand, the poor search capabilities of Lamcts-Turbo, regardless of the available budget. For higher dimensions, Fig. 5 shows that BO-based solvers become less competitive, although they still perform among the best for the smallest budgets (25 and 50 evaluations). For larger budgets, we see that CMA and DE start becoming more competitive, which is in line with [51]. However, Fig. 5f shows a supremacy of HyperOpt, which is designed for large-scale optimization for models with hundreds of parameters, and PSO, which performs surprisingly well considering that it is a classical heuristic originating from the BBO field and rarely used in ML applications. It can also be noted that Cobyla does not perform as well as for BBOB, while PSO performs constantly well for intermediate budgets from 50 to 400 total evaluations.
Fig. 6 extends the present results to budgets 800, 1600, and 3200 for both tiny and big neural nets. Here, as the total budget increases, a smaller set of algorithms are compared: AX is affordable for runs on tiny neural nets up to a budget of 800 evaluations, and BO and Lamcts-Turbo become too expensive when exceeding a total budget of 1600 evaluations for big neural nets. The RL-specific algorithms, SpecialRL, and NGOpt16RL achieve the best performance for the largest budgets, followed by the various CMA and DE versions, leaving PSO and HyperOpt behind.
@@ V Conclusion
In general, we found that the OpenAI Gym benchmark is very sensitive to variable scaling. While BBOB focuses on translations of optima and might therefore favor algorithms tuned for this setting, direct policy search for OpenAI Gym has an unbounded domain and depends differently on the initialization scaling and the ability of the algorithm to change scaling as needed. We optimized the scaling to Bayesian Optimization methods for obtaining results in Fig. 4, and then we ran results in Fig. 5 without any change.
@@ Appendix A Discussion: differences between BBOB and Nevergrad’s YABBOB
There are differences between the main benchmarking suites.
@@ Appendix C How to run the experiments on OpenAI Gym
However, in order to run a reduced experiment like the one presented in this paper, we changed the definition of the budget, dimension, and scaling (budget $25,50,100,200,400,800$, scaling-factor $1$, and add a limit $40$ to the dimension) in the ng_full_gym experiment in nevergrad/benchmarks/gymexperiments.py (Line 80) for obtaining the setup as in Section III.
@@ Appendix D Bigger neural nets for OpenAI Gym
Fig. 5 is a more complete version of the Ng-Full-Gym problem, where the neural factor parameter, which scales the size of the neural networks, is equal to 3. This does not change the optimization methods, only the scale of the problems towards a greater dimension. Dimensions up to 264 are considered. This benchmark is unbounded, the the scale of the algorithms (i.e., the standard deviation of the first samples) is therefore particularly critical and makes a fair comparison difficult.
@@ Appendix E Budgets 800, 1600, and 3200
While Figs. 4 and 5 present results restricted to budgets of at most $400$ function evaluations, Fig. 6 presents results for larger budgets of 800, 1600, and 3200 evaluations applied to multi-deterministic Open AI Gym with both tiny and (for methods which are computationally cheap enough) bigger neural nets. As in [51], CMA or NGOpt get better as the budget grows. We also find that, while most BO-based methods weaken, HyperOpt performs satisfactorily.
@@
Figure 6: Extension of Fig. 4 and 5, for budget 800 (top), 1600 (center), and 3200 (bottom) for algorithms that were sufficiently fast (faster than 3 days wall-clock time).

### v3

# Optimizing with Low Budgets: a Comparison on the Black-box Optimization Benchmarking Suite and OpenAI Gym

@@ III Benchmark Problems
Winning percentage rates are evaluated at fixed budgets of 25, 50, 100, 200, and 400 function evaluations for both settings (neural factor $1$ and $3$), while bigger budgets are also considered for the bigger networks (see Sections IV and V in the Supplementary Material).
@@ IV Results of the Empirical Comparison
Fig. 2 shows that both versions of CMA – defaultcma and cmafmin2 (and NGOpt16, which uses CMA as a component) – perform better than all other solvers when a larger budget of function evaluations is available.
@@
We add the following algorithms to our comparison: QORandomSearch [64], MetaTuneRecentering [65], and MetaRecentering [66] are variants of random search that are fully parallel and reduce redundancies compared to random search. We also add NGOpt16RL and SpecialRL from Nevergrad: NGOpt16RL is Nevergrad’s wizard NGOpt optimized specifically for RL problems; SpecialRL runs the base NGOpt16RL for half of the budget and then uses test-based population size adaptation (TBPSA) [69] in the second half. Moreover, since the OpenAI Gym suite contains problems that are unbounded with optima at a small scale, we add algorithms that can adjust the scale of the search. GeneticDE and RotatedTwoPointsDE, for example, are designed to extrapolate the scale from some variables to others and find the right scale before running the classic DE [70] algorithm. MetaTuneRecentering performs a DoE entirely designed to choose the right scale depending on the relationship between budget and dimension. MultiCMA is a tentative robustification of CMA that runs three copies of CMA. We also consider a (1+1) evolutionary algorithm [71], which is a simple and fast state-of-the-art heuristic for adjusting the scale, Diagonal CMA [26], which uses a diagonal covariance matrix, and its scaled version, Scaled Diagonal CMA [5]. Finally, we add to our portfolio Lamcts-Turbo [18], an improved version of vanilla LA-MCTS coupled with Turbo to draw new samples in the selected node of the Monte Carlo search tree at each iteration. For a complete list of all algorithms compared, see Section VI in the Supplementary Material.
@@
Figure 4: Same results as reported in Fig. 3, but aggregated comparison as provided by Nevergrad [5]. Row A column B shows the frequency at which method A outperformed method B for the given budget. 9 distinct problems per budget. We include only problems for which the dimension is $D<50$. Methods are ranked per average winning rate; best methods are listed first. Note that winning rates are all very close to each other: only PSO is significantly better. Fig. 5 presents similar experiments but with bigger neural nets. Fig. 1 in the Supplementary Material extends the present results to budgets 800, 1600, and 3200.
@@
Figure 5: Same as Fig. 4, but with bigger nets (neural factor 3 in Nevergrad’s benchmark scaling). 11 distinct problems per budget. We truncated at dimension $\leq 264$. Dimension ranges from 24 to 264 instead of 8 to 40 in Fig. 4. Due to the computational cost, it was not possible to finish the runs for SMAC. Fig. 1 in the Supplementary Material extends the present results to budgets 800, 1600, and 3200.
For an in-depth comparison, we also present aggregate plots based on average winning rates (e.g., Fig. 4), where we observe good results for the BO-based methods: BO is the most powerful solver for the lowest budget, while AX, NGOpt16RL and PSO turn out to be the best algorithms (within the confidence intervals) when aggregating results across all budgets (Fig. 4f). This means that vanilla BO is actually good for fast low-precision approximations on difficult problems, whereas AX, NGOpt16RL and PSO can be recommended when optimizing low-dimensional problems with low to medium budgets. Furthermore, Fig. 4 confirms, on the one hand, the good performance of NGOpt16RL, PSO, and HyperOpt for benchmarks with tiny neural networks, and, on the other hand, the poor search capabilities of Lamcts-Turbo, regardless of the available budget. For higher dimensions, Fig. 5 shows that BO-based solvers become less competitive, although they still perform among the best for the smallest budgets (25 and 50 evaluations). For larger budgets, we see that CMA and DE start becoming more competitive, which is in line with [49]. However, Fig. 5f shows a supremacy of HyperOpt, which is designed for large-scale optimization for models with hundreds of parameters, and PSO, which performs surprisingly well considering that it is a classical heuristic originating from the BBO field and rarely used in ML applications. It can also be noted that Cobyla does not perform as well as for BBOB, while PSO performs constantly well for intermediate budgets from 50 to 400 total evaluations.
Fig. 1 in the Supplementary Material extends the present results to budgets 800, 1600, and 3200 for both tiny and big neural nets. Here, as the total budget increases, a smaller set of algorithms are compared: AX is affordable for runs on tiny neural nets up to a budget of 800 evaluations, and BO and Lamcts-Turbo become too expensive when exceeding a total budget of 1600 evaluations for big neural nets. The RL-specific algorithms, SpecialRL, and NGOpt16RL achieve the best performance for the largest budgets, followed by the various CMA and DE versions, leaving PSO and HyperOpt behind.
@@ V Conclusion
In general, we found that the OpenAI Gym benchmark is very sensitive to variable scaling. While BBOB focuses on translations of optima and might therefore favor algorithms tuned for this setting, direct policy search for OpenAI Gym has an unbounded domain and depends differently on the initialization scaling and the ability of the algorithm to change scaling as needed. We optimized the scaling to Bayesian Optimization methods for obtaining results in Fig. 4, and then we ran the experiments to get the (possibly more neutral) results in Fig. 5 without any change.
@@ Appendix A Discussion: differences between BBOB and Nevergrad’s YABBOB
There are differences between the benchmarking suites.
@@ Appendix C How to run the experiments on OpenAI Gym
However, in order to run a reduced experiment like the one presented in this paper, we changed the definition of the budget, dimension, and scaling (budget $25,50,100,200,400,800$, scaling-factor $1$, and add a limit $40$ to the dimension) in the ng_full_gym experiment in nevergrad/benchmarks/gymexperiments.py (Line 80) for obtaining the setup as in Section III of the main paper.
@@ Appendix D Bigger neural nets for OpenAI Gym
Here, we offer additional details about the configuration employed to produce the results depicted in Fig. 5 in the main manuscript. Bigger neural networks provide a more complete version of the Ng-Full-Gym problem, where the neural factor parameter, which scales the size of the neural networks, is equal to 3. This does not change the optimization methods, only the scale of the problems towards a greater dimension. Dimensions up to 264 are considered. This benchmark is unbounded, the the scale of the algorithms (i.e., the standard deviation of the first samples) is therefore particularly critical and makes a fair comparison difficult.
@@ Appendix E Budgets 800, 1600, and 3200
While Figs. 4 and 5 in the main paper present results restricted to budgets of at most $400$ function evaluations, Fig. 6 presents results for larger budgets of 800, 1600, and 3200 evaluations applied to multi-deterministic Open AI Gym with both tiny and (for methods which are computationally cheap enough) bigger neural nets. As in [49], CMA or NGOpt get better as the budget grows. We also find that, while most BO-based methods weaken, HyperOpt performs satisfactorily.
@@
Figure 6: Extension of Figs. 4 and 5 in the main manuscript, for budget 800, 1600, and 3200, for algorithms that were sufficiently fast (faster than 3 days of wall-clock time).